OpenStudy (anonymous):

In 2000, there were about 200 million vehicles and about 279 million people in a certain country. the number of vehicles has been growing at 4% a year, while the population has been growing at 1% a year. (a) Write a formula for the number of vehicles (in millions) as a function of t, the number of years since 2000. V(t) =

OpenStudy (anonymous):

I am not quite sure about your question: do you mean that you want a formula just to show the realtionship between the number of vehicle(in M) and the time t(in years and since 2000)? if so ,the answer should be \[N _{t}==(200)*1.04^t\], where N is the number of the \[N _{t}\] is the number of vechiles in year t. Hopefully, this is what you want to know. and by the way~ you may want to ask the the average number of vechiles per people in year t if it is in this case, the answer should be the following: \[A _{t}=( N _{t})/[P _{t}\], where, \[N _{t}=(200)*1.04^t\] and \[P _{t}=279*(1+1/100)^{t}\]